# CLiMB: A Domain-Informed Novelty Detection Clustering Framework for Scientific Discovery

### v1

@@ Abstract
In data-driven scientific discovery, a challenge lies in classifying well-characterized phenomena while identifying novel anomalies. Current semi-supervised clustering algorithms do not always fully address this duality, often assuming that supervisory signals are globally representative. Consequently, methods often enforce rigid constraints that suppress unanticipated patterns or require a pre-specified number of clusters, rendering them ineffective for genuine novelty detection. To bridge this gap, we introduce CLiMB (CLustering in Multiphase Boundaries), a domain-informed framework decoupling the exploitation of prior knowledge from the exploration of unknown structures. Using a sequential two-phase approach, CLiMB first anchors known clusters using constrained partitioning, and subsequently applies density-based clustering to residual data to reveal arbitrary topologies. We demonstrate this framework on RR Lyrae stars data from the Gaia Data Release 3. CLiMB attains an Adjusted Rand Index of 0.829 with 90% seed coverage in recovering known Milky Way substructures, drastically outperforming heuristic and constraint-based baselines, which stagnate below 0.20. Furthermore, sensitivity analysis confirms CLiMB’s superior data efficiency, showing monotonic improvement as knowledge increases. Finally, the framework successfully isolates three dynamical features (Shiva, Shakti, and the Galactic Disk) in the unlabelled field, validating its potential for scientific discovery.
Keywords: novelty detection, semi-supervised clustering, constrained clustering, density-based clustering, domain knowledge integration
@@ 1 Introduction
In modern scientific discovery, data analysis frequently follows a dual objective: the precise classification of well-characterized phenomena and the simultaneous identification of novel, structurally distinct components. This challenge is ubiquitous in domains ranging from bioinformatics to astronomy, where vast datasets contain both known data instances and rare, unobserved events. While Supervised Learning excels at recovering known structures and Unsupervised Learning provides unbiased exploration, neither paradigm individually addresses this duality. Semi-Supervised Clustering (SSC) has emerged as the bridge between these extremes, aiming to improve partitioning accuracy by incorporating limited domain knowledge, typically in the form of pairwise constraints or labeled seeds (Basu and Mooney, 2005; Wagstaff et al., 2001).
However, a systematic review of the field reveals critical limitations in current SSC methodologies when applied to scientific novelty detection. The literature is predominantly driven by constraint-based methods utilizing must-link and cannot-link constraints (Basu et al., 2004b; Jiang et al., 2013). While effective for enhancing cluster purity, these methods typically operate under the assumption that the provided supervision is “representative” of the underlying data distribution (González-Almagro et al., 2023). This assumption forces algorithms to propagate constraints across the entire dataset, often distorting or suppressing the emergence of novel clusters that do not align with the priors. Furthermore, the majority of these frameworks, including variants of COP-KMeans and PC-KMeans, require the number of clusters ($k$) to be pre-specified (Wacquet et al., 2011; Xiong et al., 2014), a requirement that is fundamentally incompatible with the discovery of unanticipated structural components.
@@
The application domain further highlights this gap. While bioinformatics has seen extensive SSC application—particularly in gene expression and document clustering (Gu et al., 2013; Jiang et al., 2013), astronomical applications have received minimal explicit attention in the semi-supervised literature (Baumann and Hochbaum, 2022). This is a significant oversight, as current galactic surveys (e.g., Gaia, Gaia Collaboration et al. 2016) produce high-dimensional kinematic data in which known stellar streams coexist with undiscovered accretion remnants.
@@
The application of the CLiMB framework to the Gaia Data Release 3 (DR3, Gaia Collaboration et al. 2023) catalog of RR Lyrae stars (RRLs) described in this paper yielded new insights into the dynamical architecture of the Milky Way (MW) halo. While traditional clustering approaches struggled with the sparsity of the dataset, CLiMB’s two-phase architecture proved robust when applied to the RRL sample. We demonstrate that CLiMB significantly outperforms heuristic and constraint-based baselines, such as C-DBSCAN (Ruiz et al., 2007) and heuristic SS-DBSCAN (Abdulhameed et al., 2024), in recovering known MW substructures while effectively identifying subtle dynamical features, such as Shakti and Shiva, which remain challenging for traditional methods. In the initial constrained phase, the algorithm not only recovered eight known substructures, such as the Gaia Sausage/Enceladus (GSE; Belokurov et al. 2018; Helmi et al. 2018) and Sequoia (Myeong et al., 2019; Matsuno et al., 2019), based on integral-of-motion similarity. The most significant outcome, however, emerged during the exploratory second phase: upon removing the known structures, the algorithm successfully unveiled three previously unassigned dynamical groups in the residual data. These corresponded to the Galactic Disk and the distinct accreted substructures Shiva and Shakti (Malhan and Rix, 2024). This ability to isolate novel features from the unlabelled field resulted in the characterization of 11 total substructures.
@@ 2.2.1 Advanced Initialization and Metric Flexibility
  • Mahalanobis Distance: $d(\mathbf{x}_{i},\mathbf{x}_{j})=\sqrt{(\mathbf{x}_{i}-\mathbf{x}_{j})^{T}\mathbf{V}^{-1}(\mathbf{x}_{i}-\mathbf{x}_{j})}$, where $\mathbf{V}$ is the covariance matrix of the data.
@@ 2.4.3 Implementation Details
3 Related Works
Semi-supervised clustering (SSC) constitutes a critical subfield of machine learning that addresses the inherent ill-posedness of unsupervised clustering by incorporating limited supervisory information into the partitioning process. This supervision, typically manifested as class labels for a data subset or as pairwise constraints (i.e., must-link and cannot-link pairs), serves to regularize the problem, guiding the algorithm toward a semantically meaningful solution that aligns with available domain knowledge as described in Qin et al. (2019). The evolution of SSC methodologies reveals a trajectory from direct algorithmic modifications to more sophisticated data-space transformations and, ultimately, to hybrid multi-stage frameworks. This review charts this evolution to identify persistent challenges and contextualize the contribution of our proposed algorithm, CLiMB.
3.1 Evolution of Methodological Paradigms in SSC
The primary paradigms for integrating supervision in clustering can be broadly categorized into two major approaches: direct modification of the clustering algorithm and transformation of the underlying data representation. Early work predominantly focused on the former. One foundational strategy involves informed initialization, or seeding, where labeled data points are used to define the initial cluster centroids. By commencing the iterative optimization process from a more favorable position in the solution space, such methods mitigate the risk of convergence to suboptimal local minima, thereby enhancing both the accuracy and stability of the final partition (Basu et al., 2002).
A more flexible form of algorithmic-level integration is realized through constraint-based clustering. Foundational algorithms like Cop-Kmeans (Wagstaff et al., 2001) modify the cluster assignment step to enforce must-link and cannot-link constraints procedurally. This principle was later formalized more rigorously through the modification of the clustering objective function, as exemplified by PCKmeans (Basu et al., 2004a), and has been successfully extended to other clustering families, including hierarchical and spectral methods (Davidson and Ravi, 2005). This line of work continues to evolve, as seen in Yang et al. (2012), who introduce ICop-Kmeans, an enhanced variant that addresses constraint violations by leveraging object certainty and weighted co-association, further integrated with a novel constrained self-organizing map to refine semi-supervised clustering. This paradigm was also adapted for density-based algorithms, which are inherently better suited for discovering non-globular clusters. For instance, C-DBSCAN first partitions the data into numerous local micro-clusters while strictly respecting cannot-link constraints, and only subsequently merges these based on must-link constraints and proximity (Ruiz et al., 2007).
A distinct but related approach involves modifying algorithmic behavior not with specific pairwise constraints, but with domain-guided heuristics. Instead of providing examples, this paradigm injects general rules based on domain knowledge to guide the clustering process. The heuristic SS-DBSCAN method, inspired by the work of Abdulhameed et al. (2024), exemplifies this by altering the core point selection criteria of DBSCAN. In this model, a point can only initiate a new cluster if it is both in a dense region and satisfies an external, user-defined ”importance” function. While this allows for the integration of high-level domain knowledge, the effectiveness of such methods is highly dependent on the quality and generalizability of the handcrafted heuristic rule.
A subsequent and significant shift in SSC research moved beyond algorithmic adaptation to focus on clustering metric learning. Instead of modifying the clustering process, this paradigm aims to learn a distance metric under which the data’s geometric structure inherently conforms to the supervisory information. The seminal work by Xing et al. (2002) framed this as a convex optimization problem to learn a Mahalanobis distance metric that minimizes distances between must-link pairs while maximizing them for cannot-link pairs. The power of this approach lies in its ability to reshape the feature space topology as in Ye et al. (2007); Zhang et al. (2012), often leading to superior performance as the subsequent application of even a simple clustering algorithm can yield accurate results. This concept remains highly relevant, with modern deep clustering methods like Semi-Supervised Deep Embedded Clustering (SDEC) employing pairwise constraints to regularize the learned latent embedding space, thereby ensuring that the learned features are discriminative and well-suited for clustering (Ren et al., 2019).
The limitations of these single-phase approaches, particularly their sensitivity to noisy supervision and their rigidity in discovering previously unseen cluster structures, motivated the development of multi-step and hybrid frameworks. These more advanced methods often decouple the clustering process into distinct stages of supervision-guided partitioning and unconstrained exploration. For example, active learning frameworks refine an initial clustering by interactively querying for constraints on the most ambiguous data points (Grira et al., 2008). Other adaptive methods employ a multi-stage process wherein parameters for a density-based clustering algorithm are determined using both labeled and unlabeled data prior to the final partitioning (Yang et al., 2017). As a comprehensive survey by Cai et al. (2023) highlights, the field is replete with diverse techniques, but this trend toward multi-stage processing indicates a growing recognition of the need for more flexible and robust algorithmic designs.
3.2 Persistent Challenges and Identified Research Gaps
Despite considerable progress, the efficacy of existing SSC algorithms is predicated on several key assumptions, the violation of which exposes persistent challenges. A primary issue is the reliability of the supervisory signal. The majority of algorithms presuppose that provided labels and constraints are accurate, an assumption that is frequently untenable in real-world applications. While “safe” SSC methods have been proposed to mitigate the impact of noisy supervision, for instance by assigning confidence weights to labeled instances (Gan et al., 2019), this remains an open and critical research area.
Furthermore, a fundamental tension exists in the trade-off between exploitation and exploration, that is, between adhering to the provided supervision and discovering the intrinsic structure of the unlabeled data. Over-reliance on strong constraints can lead to a model that satisfies the supervision but fails to capture novel or emergent clusters, effectively constraining the discovery process. This exposes a critical gap in the literature: the absence of a unified framework designed explicitly to operate on datasets containing both well-characterized and entirely novel components. Current methods implicitly assume that the limited supervision is representative of the global class distribution. They lack a formal mechanism to first isolate and model the “known” data components based on supervision, and then perform unconstrained, exploratory clustering on the residual data. This limitation is particularly acute in domains such as network security, where known attack signatures must be identified alongside novel threats, or in bioinformatics, where new cell types are discovered from datasets containing previously annotated ones.
Our work, CLiMB, is designed to directly address this gap. It introduces a sequential, two-phase architecture that systematically decouples the exploitation of prior knowledge from the exploratory discovery of unknown patterns. By first partitioning the data space according to supervisory seeds and subsequently applying an unconstrained clustering algorithm to the remaining data, CLiMB provides a principled solution for clustering in multi-phase boundaries where both confirmation and discovery are required.
4 Experiments
@@
4.1 Galactic Archaeology with RRLs from Gaia DR3
@@
RRLs are pulsating variable stars, serving as ideal tracers of old, metal-poor stellar populations prevalent in the MW halo. RRLs are excellent standard candles for measuring distances due to their well-defined luminosity-metallicity relations in the visual bands and period-luminosity-metallicity relations in the near- and mid-infrared bands (e.g., Longmore et al. 1986; Clementini et al. 2003; Bono et al. 2003; Sollima et al. 2008; Muraveva et al. 2018). Additionally, RRLs are effective metallicity tracers, as their metallicity can be estimated from photometric parameters, such as the pulsation period and Fourier decomposition parameters of their light curves, without requiring spectroscopic data (e.g., Jurcsik and Kovacs 1996; Morgan et al. 2007; Muraveva et al. 2025). As remnants tracers of ancient stellar systems, RRLs are particularly valuable for studying the dynamical history of the Galaxy, including past accretion events that have shaped its halo (Massari et al., 2019; Helmi, 2020). These events, involving the merger of dwarf galaxies or globular clusters, leave kinematic and spatial signatures in the form of stellar streams and substructures, which can be traced using integrals of motion, such as the total energy ($E$), the angular momentum along the $Z$-axis ($L_{z}$), and the perpendicular component of angular momentum ($L_{\perp}$) (Helmi and Tim de Zeeuw, 2000; Ceccarelli et al., 2024; Dodd et al., 2023).
The Gaia mission (Gaia Collaboration et al., 2016), launched by the European Space Agency in 2013, has revolutionized Galactic astronomy by providing unprecedented astrometric, photometric, and spectroscopic data for over a billion stars. In both Muraveva et al. (in prep.) and the present study, we selected a sample of 4,933 RRLs, as shown in Figure 1, for which positions, proper motions, parallaxes, and radial velocities are available in Gaia DR3. We then used this six-dimensional phase-space information to compute integrals of motion and applied clustering algorithms to identify dynamical substructures in the MW halo associated with past accretion events.
From a machine learning perspective, we project this high-dimensional physical data into a compact 3-dimensional feature space $\mathbf{X}\in\mathbb{R}^{n\times 3}$. This space defines the dynamical manifold of the stellar streams and constitutes the actual input for our clustering framework:
@@
4.2 Data Preparation and Preprocessing

### v2

# CLiMB: A Domain-Informed Novelty Detection Clustering Framework for Galactic Archaeology and Scientific Discovery

@@ Abstract
In data-driven scientific discovery, a challenge lies in classifying well-characterized phenomena while identifying novel anomalies. Current semi-supervised clustering algorithms do not always fully address this duality, often assuming that supervisory signals are globally representative. Consequently, methods often enforce rigid constraints that suppress unanticipated patterns or require a pre-specified number of clusters, rendering them ineffective for genuine novelty detection. To bridge this gap, we introduce CLiMB (CLustering in Multiphase Boundaries), a domain-informed framework decoupling the exploitation of prior knowledge from the exploration of unknown structures. Using a sequential two-phase approach, CLiMB first anchors known clusters using metric-adaptive constrained partitioning, and subsequently applies density-based clustering to residual data to reveal arbitrary topologies. We demonstrate this framework on RR Lyrae stars data from the Gaia Data Release 3. CLiMB attains an Adjusted Rand Index of 0.829 with 90% seed coverage in recovering known Milky Way substructures, outperforming heuristic and constraint-based baselines, which stagnate below 0.20. Furthermore, sensitivity analysis confirms CLiMB’s superior data efficiency, showing monotonic improvement as knowledge increases. Finally, the framework successfully isolates three distinct dynamical features (Shiva, Shakti, and the Galactic Disk) in the unlabelled field, validating its potential for scientific discovery.
keywords:
Machine learning , Novelty detection , Semi-supervised clustering , Domain knowledge integration , Kinematics and dynamics , RR Lyrae , Data analysis
journal: Astronomy &\& Computing
[…]
[inst1]organization=INAF, Osservatorio di Astrofisica e Scienza dello Spazio, addressline=via Piero Gobetti 93/3, city=Bologna, postcode=40129, country=Italy
[…]
[inst2]organization=TOELT LLC, Machine Learning Research and Development Department, city=Winterthur, postcode=8406, state=Zurich, country=Switzerland
[…]
[inst3]organization=Computer Science Department, Lucerne University of Applied Sciences and Arts, city=Luzern, postcode=6002, country=Switzerland
@@ 1 Introduction
[…]
RR Lyrae stars (RRLs) are pulsating variable stars, serving as ideal tracers of old ($\geq 10$ Gyr), metal-poor stellar populations prevalent in the MW halo. RRLs are excellent standard candles for measuring distances due to their well-defined luminosity-metallicity relations in the visual bands and period-luminosity-metallicity relations in the near- and mid-infrared bands (e.g., Longmore et al. 1986; Clementini et al. 2003; Bono et al. 2003; Sollima et al. 2008; Muraveva et al. 2018). Additionally, RRLs are effective metallicity tracers, as their metallicity can be estimated from photometric parameters, such as the pulsation period and Fourier decomposition parameters of their light curves, without requiring spectroscopic data (e.g., Jurcsik and Kovacs 1996; Morgan et al. 2007; Muraveva et al. 2025).
As remnants tracers of ancient stellar systems, RRLs are particularly valuable for studying the dynamical history of the Galaxy, including past accretion events that have shaped its halo (Massari et al., 2019; Helmi, 2020). These events, involving the merger of dwarf galaxies or globular clusters, leave kinematic and spatial signatures in the form of stellar streams and substructures, which can be traced using IoM, such as the total energy ($E$), the angular momentum along the $Z$-axis ($L_{z}$), and the perpendicular component of angular momentum ($L_{\perp}$) (Helmi and Tim de Zeeuw, 2000; Ceccarelli et al., 2024; Dodd et al., 2023). The ESA Gaia mission (Gaia Collaboration et al., 2016) has revolutionized this field, providing high-precision astrometric and spectroscopic data. In this study, we leverage a sample of 4,933 RRLs from Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023), for which full 6D phase-space information (coordinates, proper motions, distances, and radial velocities) is available, to compute IoM and evaluate our clustering framework.
Traditional unsupervised clustering approaches treat all data points as equally unknown, ignoring the vast amount of domain knowledge already accumulated about the Galactic halo. Conversely, supervised classification methods are limited to assigning stars to known groups, inherently failing to discover new accretion events. This tension underscores the importance of developing a framework in modern Galactic archaeology that can precisely recover well-characterized substructures while remaining sensitive to novel, unmapped features.
@@
The application domain further highlights this gap. While bioinformatics has seen extensive semi-supervised clustering (SSC) application—particularly in gene expression and document clustering (Gu et al., 2013; Jiang et al., 2013), astronomical applications have received minimal explicit attention in the semi-supervised literature (Baumann and Hochbaum, 2022). This is a significant oversight, as current galactic surveys (e.g., Gaia, Gaia Collaboration et al. 2016) produce high-dimensional kinematic data in which known stellar streams coexist with undiscovered accretion remnants.
@@
We applied the CLiMB framework to the Gaia DR3 catalog of RRLs to dissect the dynamical architecture of the Milky Way (MW) halo. The framework’s two-phase architecture proved robust against the dataset’s sparsity, significantly outperforming heuristic baselines like C-DBSCAN (Ruiz et al., 2007) and SS-DBSCAN (Abdulhameed et al., 2024). In the initial constrained phase, CLiMB successfully recovered eight known substructures based on IoM similarity, including the Gaia Sausage/Enceladus (GSE) and Sequoia. Crucially, the exploratory second phase unveiled three previously unassigned dynamical groups in the residual data: the Galactic Disk and the distinct accreted substructures Shiva and Shakti (Malhan and Rix, 2024). By isolating these novel features from the unlabelled field, features that remain elusive for traditional clustering methods, CLiMB characterized a total of 11 substructures, offering new insights into the MW’s assembly history.
The paper is organised as follows: Section 2 reviews the relevant literature in semi-supervised clustering and its applications in Galactic archaeology. Section 3 details the mathematical architecture of the CLiMB framework and its two-phase clustering approach. Section 4 describes the experimental setup, including the Gaia RRL dataset preparation, the calculation of IoM, and the benchmarking protocol against baseline algorithms. Section 5 presents the comparative analysis of the results, the sensitivity study on prior knowledge, and the scientific validation of the discovered substructures. Finally, Section 6 provides a summary of our findings and discusses future directions for novelty detection in large-scale astronomical surveys.
2 Related Works
Semi-supervised clustering (SSC) constitutes a critical subfield of machine learning that addresses the inherent ill-posedness of unsupervised clustering by incorporating limited supervisory information into the partitioning process. This supervision, typically manifested as class labels for a data subset or as pairwise constraints (i.e., must-link and cannot-link pairs), serves to regularize the problem, guiding the algorithm toward a semantically meaningful solution that aligns with available domain knowledge as described in (Basu and Mooney, 2005; Qin et al., 2019).
2.1 Paradigms in Semi-Supervised Clustering
The primary paradigms for integrating supervision in clustering can be broadly categorized into two major approaches: the direct modification of the clustering algorithm and the transformation of the underlying data representation. Early work predominantly focused on algorithm-level integration through informed initialization or seeding. In this paradigm, labeled data points are used to define the initial cluster centroids. By commencing the iterative optimization process from a more favorable position in the solution space, such methods mitigate the risk of convergence to suboptimal local minima, enhancing both the accuracy and stability of the final partition (Basu et al., 2002). This principle is a direct precursor to the initial phase of our proposed framework, although traditional seeding lacks the spatial boundaries necessary for novelty detection.
A more flexible form of algorithmic-level integration is realized through constraint-based clustering. Foundational algorithms like COP-Kmeans (Wagstaff et al., 2001) and PCKmeans (Basu et al., 2004) modify the cluster assignment step to enforce must-link and cannot-link constraints. This paradigm was later adapted for density-based algorithms, which are inherently better suited for discovering non-globular clusters. For instance, C-DBSCAN (Ruiz et al., 2007) first partitions the data into numerous local micro-clusters while strictly respecting cannot-link constraints, and only subsequently merges them based on must-link constraints and proximity. While robust in enforcing local labels, C-DBSCAN is prone to ”chain-reaction” merging in high-density regions, where distinct stellar streams might be incorrectly connected by bridges of background noise.
A distinct approach involves heuristic-guided clustering, which modifies algorithmic behavior not with specific pairwise constraints, but with domain-guided rules. Instead of providing examples, this paradigm injects general rules based on domain knowledge to guide the clustering process. The SS-DBSCAN method, as described in Abdulhameed et al. (2024), exemplifies this by altering the core point selection criteria of DBSCAN. In this model, a point can only initiate a new cluster if it is both in a dense region and satisfies an external, user-defined ”importance” function. While this allows for the integration of high-level domain knowledge, the effectiveness of such methods is highly dependent on the quality and generalizability of the handcrafted heuristic rule.
Moreover, metric learning aims to learn a distance function—often a Mahalanobis metric—under which the data’s geometric structure conforms to the supervisory information (Xing et al., 2002; Ye et al., 2007). While powerful in reshaping the feature space, metric learning often struggles with ”local validity”; a distance metric optimized for a specific subset of known classes may be ill-suited for the geometry of unknown regions.
More recently, the field has moved toward hybrid multi-stage frameworks and deep clustering (Ren et al., 2019). These advanced methods often attempt to decouple the clustering process into distinct stages of supervision-guided partitioning and unconstrained exploration. Yet, as noted by Ienco and Interdonato (2022), the tension between exploitation (adhering to prior knowledge) and exploration (discovering intrinsic structure) remains a persistent challenge. CLiMB is designed to resolve this tension by formally decoupling these two objectives into a sequential, multiphase architecture.
2.2 Clustering in Galactic Archaeology
In the context of Galactic archaeology, the identification of stellar streams has traditionally relied on density-based unsupervised algorithms, most notably DBSCAN (Ester et al., 1996) and HDBSCAN (Campello et al., 2013), due to their ability to detect non-convex shapes and handle background noise. These methods have been instrumental in mapping the MW’s accretion history (Koppelman et al., 2019; Cabrera Garcia et al., 2024). However, purely unsupervised methods are highly sensitive to hyperparameter tuning and often struggle with the ”varying density” problem inherent in Galactic surveys, where dense disk populations coexist with sparse halo streams.
To incorporate physical priors, model-dependent algorithms like STREAMFINDER (Malhan and Ibata, 2018) utilize stellar evolutionary models and orbital integration to find coherent structures. While physically rigorous, these methods are computationally intensive and potentially biased toward structures that fit the assumed potential. Conversely, purely data-driven approaches like StarGO (Yuan et al., 2018) utilize Self-Organizing Maps (SOMs) to find over-densities in high-dimensional space but lack a formal mechanism to integrate existing catalogs of known structures. Furthermore, while these methods focus on mapping the global manifold, they often lack the fine-grained control needed to isolate specific overlapping substructures during the learning process.
2.3 Bridging the Gap: Decoupling and Novelty Detection
The challenge of performing novelty detection within a semi-supervised framework remains an open research area. Most density-based SSC variants, such as SS-DBSCAN (Lelis and Sander, 2009) or the heuristic-guided SS-DBSCAN (Abdulhameed et al., 2024), attempt to guide cluster formation through density-connected components or domain rules. However, in high-density regions where different dynamical entities overlap, label propagation can lead to ”chaining effects,” where known labels bleed into unrelated structures.
Recent hybrid frameworks have proposed multi-stage architectures to address these rigidities. For instance, the $S^{3}C^{2}$ algorithm decomposes clustering into binary classification and supervised grouping (Śmieja et al., 2020), while other methods utilize multiresolution autoencoders to capture different structural granularities (Ienco and Pensa, 2018). CLiMB builds upon the observation that separating constraint satisfaction (exploitation) from density-based exploration is crucial for scientific discovery (Ienco and Interdonato, 2022). By utilizing a two-phase architecture, CLiMB avoids the pitfalls of global constraint propagation, ensuring that prior knowledge stabilizes known structures without constraining the algorithm’s ability to discover novel, arbitrary topologies in the residual data.
@@ 3.2.1 Advanced Initialization and Metric Flexibility
  • 2.
    Mahalanobis Distance: $d(\mathbf{x}_{i},\mathbf{x}_{j})=\sqrt{(\mathbf{x}_{i}-\mathbf{x}_{j})^{T}\mathbf{V}^{-1}(\mathbf{x}_{i}-\mathbf{x}_{j})}$, where $\mathbf{V}$ is the covariance matrix of the data. The use of the Mahalanobis metric is particularly critical when dealing with stellar substructures in the space of integrals of motion. Unlike the Euclidean metric, which assumes isotropic (spherical) cluster shapes, the Mahalanobis distance accounts for the correlations and different scales between features such as energy ($E$) and angular momentum ($L_{z}$). In Galactic archaeology, accreted structures often appear as elongated or tilted ”shards” rather than globular distributions due to the physical constraints of tidal stripping and orbital evolution. By incorporating the covariance matrix, CLiMB can define ”ellipsoidal” boundaries that better conform to the natural dynamical manifold of stellar streams, preventing the misclassification of field stars that might be geometrically close in Euclidean space but are dynamically distinct.
  • 3.
@@ 3.4.3 Implementation Details
4 Experimental Evaluation
@@
4.1 Data Selection: RRLs from Gaia DR3
@@
Gaia DR3 provides a clean catalogue of 270,891 RRLs, processed via the Specific Object Study pipeline (SOS Cep&RRL; Clementini et al. 2023). In Muraveva et al. (2025), we refined this selection to a sample of 258,696 stars, which we consider bona fide RRLs. From this census, we selected stars with available coordinates, radial velocities, and proper motions in the Gaia Early Data Release 3 (EDR3, Gaia Collaboration et al. 2021) and DR3 (Gaia Collaboration et al., 2023) catalogues. Geometric distances were derived from Gaia EDR3 parallaxes using a Bayesian approach (Bailer-Jones et al., 2021). This selection resulted in a final sample of 4,933 RRLs (see Figure 1). For these stars, we calculated the IoM ($E,L_{z},L_{\perp}$) using the AGAMA software package (Vasiliev, 2019), adopting the MW potential from McMillan (2017).
From a machine learning perspective, we project this data into a compact 3-dimensional feature space $\mathbf{X}\in\mathbb{R}^{n\times 3}$:
@@
4.2 Preprocessing and Feature Engineering